# Revisit, Extend, and Enhance Hessian-Free Influence Functions

### v2

@@ Abstract
Influence functions serve as crucial tools for assessing sample influence. By employing the first-order Taylor extension, sample influence can be estimated without the need for expensive model retraining. However, applying influence functions directly to deep models presents challenges, primarily due to the non-convex nature of the loss function and the large size of model parameters. This difficulty not only makes computing the inverse of the Hessian matrix costly but also renders it non-existent in some cases. Various approaches, including matrix decomposition, have been explored to expedite and approximate the inversion of the Hessian matrix, with the aim of making influence functions applicable to deep models. In this paper, we revisit a specific, albeit naive, yet effective approximation method known as TracIn, and simplify it further, introducing the name Inner Product (IP). This method substitutes the inverse of the Hessian matrix with an identity matrix. We offer deeper insights into why this straightforward approximation method is effective. Furthermore, we extend its applications beyond measuring model utility to include considerations of fairness and robustness. Finally, we enhance IP through an ensemble strategy. To validate its effectiveness, we conduct experiments on synthetic data and extensive evaluations on noisy label detection, sample selection for large language model fine-tuning, and defense against adversarial attacks.
@@ 1 Introduction
Data-centric learning is a growing research field that focuses on enhancing machine learning model performance by refining the quality and characteristics of training data [42]. In contrast to model-centric approaches, which prioritize improving algorithms or optimization techniques without altering the dataset, data-centric learning involves adjusting the dataset itself—through methods like trimming, relabeling, and reweighting—while keeping the learning algorithm fixed. This approach plays a vital role in areas such as model interpretability, selecting training subsets, generating synthetic data, detecting noisy labels, improving active learning, and promoting fairness learning [11, 31].
@@
Sample influence estimation, as the foundation of data-centric learning, can be generally categorized into two categories [20]. (a) Retraining-based methods assess the sample influence by retraining the model with and without a specific sample and checking the performance change, which include the classical leave-one-out influence approach [13] and Shapley value approaches [18, 24, 30, 23]. (b) Gradient-based methods estimate influence without expensive overheads of retraining, known as influence functions. The seminal work in this category is that of [26], which utilizes a Taylor-series approximation and LiSSA optimization [1] to compute sample influences. However, the limiting assumption is that the model and loss function are convex. Despite debates on the necessity of convexity [2, 19, 4, 17], challenges persist when directly applying gradient-based methods to large models. The size of model parameters complicates calculations, particularly in obtaining the inverse of the Hessian matrix. Efforts, including matrix decomposition techniques [26, 19, 31], aim to expedite and approximate Hessian matrix inversion, therefore enhance the feasibility of influence functions for deep learning models.
Contributions. In this paper, we focus on the influence function category and revisit a specific naive yet aggressive approximation method, TracIn [46]. This method substitutes the inverse of the Hessian matrix with an identity matrix, representing it as a Hessian-free influence function—the inner product of the gradient of the validation set and a sample to be assessed, which we refer to as Inner Product (IP). We summarize our major contributions as follows:
@@
We simplify the checkpoints of TracIn, define it into IP formulation, and delve into the rationale behind this simple approximation, offering insights into why it performs well in practice.
@@
Expanding our IP framework, we extend its applicability beyond measuring sample influence on model utility to encompass considerations of fairness and robustness.
@@
To improve generalization, we introduce IP Ensemble, an innovative approach that employs dropout mechanisms to simulate a diverse array of models. IP Ensemble consolidates IP scores from these different models, thereby strengthening the method’s overall generalization capabilities.
@@
By assessing the sample influence, model performance can be enhanced by trimming detrimental samples from the training or fine-tuning set, downweighting them, or even relabeling them. Specifically, we validate the effectiveness of IP through synthetic data experiments and conduct extensive real-world evaluations using IP Ensemble. These experiments span various applications, including noisy label correction for vision data, data curation aimed at fine-tuning fairer NLP models, and defense strategies against adaptive evasion adversaries.
@@ 2 Related Work
Efficient Influence Estimation. Influence functions serve as crucial tools for estimating the individual valuation of data without requiring model retraining. However, the computation of the inverse of the Hessian matrix poses challenges for large-scale data and models. To address this issue, various approaches have been proposed to simplify or estimate the inverse of the Hessian matrix effectively. A seminal work is that of [26], which employs a Taylor-series approximation and LiSSA optimization [1] to compute sample influences. Arnoldi [48] employs the random projection and simplified Hessian matrix for acceleration. EKFAC [19] enhances Kronecker-Factored eigendecomposition for a precise Hessian approximation. More recently, DataInf [31] efficiently computes influence even for large models by replacing the inverse Hessian computation with a readily computable closed-form expression, although their framework may suffer from significant theoretical errors. TracInc [46], a straightforward yet aggressive approximation, substitutes the inverse of the Hessian matrix with an identity matrix, essentially considering gradients directly as a measure of influence. Beyond the conventional influence function that gauges sample influence on the validation set, self-influence [5, 52] computes influence using the training set alone. Moving beyond using a single model checkpoint, GEX [25] leverages a geometric ensemble of multiple checkpoints to approximate influence functions, alleviating the bilinear constraint and non-linear losses. Moreover, TDA [3] also introduces a checkpoint-based segmentation approach, which further enhances the training process by combining the techniques of implicit differentiation and unrolling. This is achieved through the utilization of EKFAC, providing a more efficient and scalable solution for complex models [19].
@@
Miscellaneous. Several studies have examined the fragility of influence functions in explaining deep learning model predictions. [2] discovers that while influence estimates may not perfectly align with leave-one-out retraining, they approximate the proximal Bregman response function, offering valuable insights for identifying influential or mislabeled examples. [4] demonstrates that the effectiveness of influence functions in neural networks varies with network architecture, depth, width, parameterization, and regularization, underscoring their fragility in deep learning due to non-convex loss functions. [17] suggests that the instability of current validation procedures, rather than non-convexity or lack of regularization, may be responsible for their unreliability. [27] expands influence functions from estimating the effects of removing one point to large groups of training samples; [38] enhance influence estimation in large-scale models by concentrating on target parameters and addressing computational instability with a robust inverse-Hessian-vector product approximation; [9] extend traditional influence functions to monitor the impact of pre-training data on fine-tuned model predictions, facilitating the identification of crucial pre-training examples.
@@ 3 Methods
Revisit and Simplify. Given a training set and a classifier with empirical risk minimization by a convex loss function , the optimal parameters of the classifier can be obtained by . To measure the influence of an individual data sample, we can train the model with and without the specific sample and see the performance change. However, the retrain-based approach is expensive for large-scale data and models. To avoid model retraining, influence functions estimate the effect of changing an infinitesimal weight of samples on a validation set , based on an impact function evaluating the quantity of interest. Considering the sample impact on model utility, i.e., the loss on the validation set, by removing this sample from the original training set, the sample influence can be estimated as follows [26]:
@@
Influence functions encounter a challenge in direct application to deep models, primarily due to the non-convex nature of the loss function and the considerable size of model parameters. This obstacle not only renders the calculation of the inverse of the Hessian matrix costly but also leads to its non-existence. Various attempts, including matrix decomposition methods [26, 19, 31], have been undertaken to expedite and approximate the inversion of the Hessian matrix, aiming to render influence functions viable for deep models. In this paper, we revisit a particular naive yet aggressive approximation method TracIn [46] by substituting the inverse of the Hessian matrix with an identity matrix, outlined as follows:
The above equation is the inner product of and the sample gradient ; therefore, we call this method Inner Product (IP). Note that TracIn incorporates multiple checkpoints to record model parameters throughout the optimization process, whereas IP only takes into account the final or converged model. We believe that sample influence should be assessed using a fixed model; simply summing sample influences at different stages fails to capture the dynamics of model optimization. The converged model can significantly diverge from its earlier stages, making the sample influences derived from initial checkpoints less accurate.
In the following, we link the IP to influence functions and provide our insights of IP on why such a naive approximation works well in practice. Let , then influence functions in Eqs. (1) and (2) can be reformulated as and . We illustrate the directions of and in Figure 1, with the influence scores roughly representing the angle between the sample gradient and or . Remarkably, although influence scores by Eqs. (1) and (2) may differ, they exhibit order-consistency in several scenarios. For instance, if two samples and have gradients in Region II or IV, then implies . However, this order-consistency does not hold in Regions I and III. Fortunately, samples in these regions are jointly recognized by and as either beneficial or detrimental, resulting in minimal practical differences. A full analysis is in Appendix A.
In general, IP emerges as a straightforward, efficient, and remarkably effective alternative to influence functions. In non-convex scenarios, approximations of the inverse Hessian matrix may introduce significant errors, rendering ineffective, even incorrect. We substantiate this observation empirically in Section 4. Conversely, while in IP may not precisely align with the optimal direction for enhancing model performance, it remains a reliable indicator for distinguishing between detrimental and beneficial samples, even for non-convex models. Intuitively, if a sample gradient aligns with the gradient from the validation set, it suggests that incorporating this sample contributes to enhancing the model’s utility. Similarly, the negative samples works oppositely.
@@
Extension. Beyond measuring the sample influence on model utility, we extend IP to assess the sample influence on fairness and robustness by modifying the impact function .
@@
Similarly, we can also measure the sample influence on adversarial robustness within the IP framework. To achieve this, we follow [11] and consider a white-box adversary [39] specific to linear models, which can be easily extended to other models and settings. To craft an adversarial sample, we take each sample in the validation set and only perturb and keep unchanged, where are the linear model coefficients, is the intercept, and controls the amount of perturbation added. In this manner, we can obtain an adversarial validation set which consists of for each sample of . Now, we can compute adversarial robustness influence for each training sample as follows:
Enhancement. The simplicity of IP offers opportunities to enhance the generalization of influence functions. In convex cases, the model parameter is both optimal and unique. However, in non-convex scenarios, the presence of local minima introduces instability and non-uniqueness into the solution. Typically, ensemble strategies are employed to bolster model generalization [14, 32]. Yet, while employing different models can enhance performance, it also escalates the costs associated with model training and complicates the calculation of influence functions. This arises from the variability in model parameters, necessitating multiple computations of the inverse of each individual Hessian matrix. The introduction of Hessian-free IP circumvents this issue, eliminating the need for costly calculations of Hessian matrices and their inverses. Drawing inspiration from dropout mechanisms, diverse models can be swiftly generated without necessitating model retraining. By computing sample gradients from various models, we propose IP Ensemble that amalgamates IP scores from distinct models. Experiments detailed in Section 5 illustrate the superior performance of the inner product ensemble over other influence function-based methods.
@@ 4 Correctness Verification on Synthetic Data
Here we verify the correctness of our Inner Product (IP) as an influence score surrogate on two synthetic datasets with convex/non-convex models. First, we use Logistic Regression on a linear dataset to demonstrate the effectiveness of IP in a convex optimization scenario. Second, we use a
Multi-Layer Perceptron on a half-moon dataset to validate the performance of IP in a non-convex optimization scenario.111The detailed characteristics of datasets and models used in this paper can be found in Appendix B. In Figure 2, A and B illustrate the training and test sets of a linearly separable dataset, consisting of 150 and 100 samples, respectively. Notably, the training set contains 10 manually generated noisy samples with incorrect labels. Similarly, D-F illustrate a non-linear separable half-moons dataset, consisting of 200 training samples including 20 noisy samples with incorrect labels, and 100 test samples. C and F display the influence score and inner product score for each training sample, calculated according to Eqs. (1) and (2), respectively. Additionally, fairness and robustness analysis is in Appendix C.1.
In the linear case depicted in Figure 2C, the inner product score serves as a reliable surrogate to distinguish detrimental samples from beneficial ones. It exhibits an almost perfect correlation and order-consistency with the influence score. Specifically, detrimental samples yield negative scores for both inner product and influence, while other samples typically show positive or nearly zero values. Figure 3 illustrates the relationship between directions among , , and the gradients of detrimental samples. It is worth noting that the angle between and is only , indicating that the identity matrix is a simple and effective surrogate of in the linear case.
However, the limitations of influence scores become apparent in the context of non-convex models, as illustrated in Figure 2F. Here, the influence scores of detrimental samples are intermingled with those of normal ones, where is not the optimal direction due to the inaccuracies in approximating the Hessian matrix. Fortunately, the IP score effectively isolates detrimental samples from inliers. Even for this non-linear dataset, remains useful for discerning detrimental samples from beneficial ones, as all detrimental samples exhibit obtuse angles with .

### v4

@@ Abstract
Influence functions serve as crucial tools for assessing sample influence. By employing the first-order Taylor expansion, sample influence can be estimated without the need for expensive model retraining. However, applying influence functions directly to deep models presents challenges, primarily due to the non-convex nature of the loss function and the large size of model parameters. This difficulty not only makes computing the inverse of the Hessian matrix costly but also renders it non-existent in some cases. In this paper, we revisit a Hessian-free method, which substitutes the inverse of the Hessian matrix with an identity matrix, and offer deeper insights into why this straightforward approximation method is effective. Furthermore, we extend its applications beyond measuring model utility to include considerations of fairness and robustness. Finally, we enhance this method through an ensemble strategy. To validate its effectiveness, we conduct experiments on synthetic data and extensive evaluations on noisy label detection, sample selection for large language model fine-tuning, and defense against adversarial attacks.
@@ 1 Introduction
Data-centric learning is a growing research field that focuses on enhancing machine learning model performance by refining the quality and characteristics of training data [43]. In contrast to model-centric approaches, which prioritize improving algorithms or optimization techniques, data-centric learning involves adjusting the dataset—through trimming, relabeling, and reweighting—while keeping the learning algorithm fixed. This approach plays a vital role in areas such as model interpretability, selecting training subsets, generating synthetic data, detecting noisy labels, improving active learning, and promoting fairness [11, 31].
@@
Sample influence estimation, as the foundation of data-centric learning, can be generally categorized into two categories [21]. (a) Retraining-based methods assess the sample influence by retraining the model with and without a specific sample and checking the performance change, which include the classical leave-one-out influence approach [13] and Shapley value approaches [19, 25, 32, 24]. (b) Gradient-based methods estimate influence without expensive overheads of retraining, known as influence functions. The seminal work in this category is that of [27], which utilizes a Taylor-series approximation and LiSSA optimization [1] to compute sample influences with the inverse of the Hessian matrix and sample gradient. However, the limiting assumption is that the model and loss function are convex. Despite debates on the necessity of convexity [3, 20, 4, 18], challenges persist when directly applying gradient-based methods to large models. The size of model parameters complicates calculations, particularly in obtaining the inverse of the Hessian matrix. Efforts, including matrix decomposition techniques [27, 20, 31], aim to expedite and approximate Hessian matrix inversion, therefore enhance the feasibility of influence functions for deep models.
Contributions. In this paper, we focus on the influence function category and revisit a specific naive yet aggressive approximation method, TracIn [47] or an early version [8]. This method substitutes the inverse of the Hessian matrix with an identity matrix, representing it as a Hessian-free influence function—the inner product of the gradient of the validation set and a sample to be assessed, which we refer to as Inner Product (IP). Rather than proposing a novel algorithm, our paper delivers an important message to the data-centric community: existing Hessian-free, simple methods, such as IP, might outperform more complex approaches that rely on dedicated Hessian inverse approximations. And we further explain the underlying reason. This finding emphasizes the importance of simplicity, efficiency, and scalability in influence estimation methods. Furthermore, it highlights the potential of these straightforward techniques to serve as a robust foundation for extending a wide range of data-centric diverse tasks. We summarize our major contributions as follows:
@@
We revisit the Hessian-free influence function, define it into IP formulation, and delve into the rationale behind this simple approximation, offering insights into why it performs well in practice.
@@
Expanding our IP framework, we extend its applicability beyond measuring sample influence on model utility to fairness and robustness.
@@
To enhance the generalization, we propose IP Ensemble, a novel approach leveraging dropout mechanisms to simulate diverse models. IP Ensemble amalgamates IP scores from these varied models, thus increasing the method’s generalization capabilities.
@@
We validate the effectiveness of IP through synthetic data experiments and conduct extensive real-world evaluations using IP Ensemble. These experiments span various applications, including noisy label correction for vision data, data curation aimed at fine-tuning fairer NLP models, and defense strategies against adaptive evasion adversaries.
@@ 2 Related Work
Efficient Influence Estimation. Influence functions serve as crucial tools for estimating the individual valuation of data without requiring model retraining. However, the computation of the inverse of the Hessian matrix poses challenges for large-scale data and models. To address this issue, various approaches have been proposed to simplify or estimate the inverse of the Hessian matrix effectively. A seminal work is that of [27], which employs a Taylor-series approximation and LiSSA optimization [1] to compute sample influences. Arnoldi [49] employs the random projection and simplified Hessian matrix for acceleration. EKFAC [20] enhances Kronecker-Factored eigendecomposition for a precise Hessian approximation. More recently, DataInf [31] efficiently computes influence even for large models by replacing the inverse Hessian computation with a readily computable closed-form expression, although their framework may suffer from significant theoretical errors. TracInc [47], a straightforward yet aggressive approximation, substitutes the inverse of the Hessian matrix with an identity matrix, essentially considering gradients directly as a measure of influence. Beyond the conventional influence function that gauges sample influence on the validation set, self-influence [5, 53] computes influence using the training set alone. Moving beyond using a single model checkpoint, GEX [26] leverages a geometric ensemble of multiple checkpoints to approximate influence functions, alleviating the bilinear constraint and non-linear losses. Moreover, TDA [2] also introduces a checkpoint-based segmentation approach, combining implicit differentiation and unrolling by using EKFAC [20]. One concurrent work [14] employs ensemble strategy to improve the efficiency of TRAK [46], a random project-based method to tackle high-dimensionaly to the Hessian matrix.
@@
Miscellaneous. Several studies have examined the fragility of influence functions in explaining deep learning model predictions. Koh et al. [28] expands influence functions from estimating the effects of removing one point to large groups of training samples. Chen et al. [9] extend traditional influence functions to monitor the impact of pre-training data on fine-tuned model predictions, facilitating the identification of crucial examples. Basu et al. [4] demonstrates that the effectiveness of influence functions in neural networks varies with network architecture, depth, width, parameterization, and regularization, underscoring their fragility in deep learning due to non-convex loss functions. Bae et al. [3] discovers that while influence estimates may not perfectly align with leave-one-out retraining, they approximate the proximal Bregman response function, offering valuable insights for identifying influential or mislabeled examples. Epifano et al. [18] suggests that the instability of current validation procedures, rather than non-convexity or lack of regularization, may be responsible for their unreliability. [39] enhance influence estimation in large-scale models by concentrating on target parameters and addressing computational instability with a robust inverse-Hessian-vector product approximation.
@@ 3 Methods
Revisit and Simplify. Given a training set and a classifier with empirical risk minimization by a convex loss function , the optimal parameters of the classifier can be obtained by . To measure the influence of an individual data sample, we can train the model with and without the specific sample and see the performance change. However, the retrain-based approach is expensive for large-scale data and models. To avoid model retraining, influence functions estimate the effect of changing an infinitesimal weight of samples on a validation set , based on an impact function evaluating the quantity of interest. Considering the sample impact on model utility, i.e., the loss on the validation set, by removing one sample from the training set, this sample influence can be estimated as follows [27]:
@@
Influence functions encounter a challenge for deep models, primarily due to the non-convex nature of the loss function and the considerable size of model parameters. This obstacle not only renders the calculation of the inverse of the Hessian matrix costly but also leads to its non-existence. Various attempts, including matrix decomposition methods [27, 20, 31], have been undertaken to expedite and approximate the inversion of the Hessian matrix, aiming to render influence functions viable for deep models. In this paper, we revisit a particular naive yet aggressive approximation method TracIn [47] or an early version [8] by substituting the inverse of the Hessian matrix with an identity matrix, outlined as follows:
The above calculation can be regarded as the inner product of two gradients; therefore, we call this method Inner Product (IP). Note that TracIn [47] incorporates multiple checkpoints to record model parameters throughout the optimization process, whereas Charpiat et al. [8] or IP only takes into account the final or converged model. While it seems appealing to consider the dynamic sample influence change according to the optimized model parameters, simply summing these simple influence scores along different checkpoints fails to account for the evolving nature of influence, potentially neutralizing conflicting values and reducing overall effectiveness, which can be verified in our experimental results (See Table 1).
In the following, we link the IP to influence functions and provide our insights of IP on why such a naive approximation works well in practice from the angle of both the traditional influence function and IP. Traditional influence functions and their approximations require the inverse of the Hessian matrix. However, the non-linear nature of data and models not only makes this computation challenging but may also lead to the non-existence of a Hessian inverse. To address this, a modified Hessian matrix is often employed, with approximations [27, 20, 31] from matrix theory used to calculate the influence score. However, choosing the regularization parameter presents a tradeoff: a small may not guarantee the existence of the inverse, while a large makes the inverse approximate the identity matrix. From the perspective of IP, it measures the similarity between two samples—the overall validation set and a specific training sample. A high IP score indicates that the target training sample is similar to the validation set, suggesting that this sample contributes positively to improving the model’s performance for both convex and non-convex models. However, in modern deep networks the Hessian is typically large, ill-conditioned, and often indefinite, so may not exist in a strict sense and must be heavily regularized. As a result, Hessian-inverse approximations can be numerically unstable, and their approximation error is difficult to quantify, especially in non-convex settings. In this work we therefore adopt the Hessian-free inner product in Eq. (2): rather than aiming to faithfully approximate , we focus on a simple and scalable gradient-alignment signal between training samples and a validation set. While IP may not always recover the exact influence value, it provides a reliable indication of the polarity of influence (beneficial vs. detrimental), which is often the main quantity needed in data-centric applications.
@@
Extension. Beyond measuring sample influence on model utility, we extend IP to assess the sample influence on fairness and robustness by modifying the impact function .
@@
Similarly, we can also measure the sample influence on adversarial robustness within the IP framework. To achieve this, we follow Chhabra et al. [11] and consider a white-box adversary [40] specific to linear models, which can be easily extended to other models and settings. To craft an adversarial sample, we take each sample in the validation set and only perturb and keep unchanged, where are the linear model coefficients, is the intercept, and controls the amount of perturbation added. In this manner, we can obtain an adversarial validation set which consists of for each sample of . Now, we can compute adversarial robustness influence for each training sample as follows:
Enhancement. The simplicity of IP offers opportunities to enhance the generalization of influence functions. In convex cases, the model parameter is both optimal and unique. However, in non-convex scenarios, the presence of local minima introduces instability and non-uniqueness into the solution. Typically, ensemble strategies are employed to bolster model generalization [15, 33]. Yet, while employing different models can enhance performance, it also escalates the costs associated with model training and complicates the calculation of influence functions. This arises from the variability in model parameters, necessitating multiple computations of the inverse of each individual Hessian matrix. The introduction of Hessian-free IP circumvents this issue, eliminating the need for costly calculations of Hessian matrices and their inverses. Drawing inspiration from dropout mechanisms, diverse models can be swiftly generated without necessitating model retraining. By computing sample gradients from various models, we propose IP Ensemble that amalgamates IP scores from distinct models. Experiments detailed in Section 5 illustrate superior performance of IP Ensemble over other influence function-based methods.
When can IP fail? Despite the above insights and enhancements, IP is not universally reliable. IP can be viewed as a Hessian-free approximation of classical influence functions, where the inverse Hessian in Eq. (1) is replaced by the identity matrix, yielding Eq. (2). Under Eq. (2), IP essentially measures the alignment between a training sample’s gradient and the validation loss gradient (i.e., the aggregated gradient over the validation set). It is worthy noting that influence functions are not accurate tools for data valuation in the deep models. Fortunately, it holds for deep models that adding more validation-like training samples tends to reduce the validation loss, which typically holds when the model is sufficiently trained and the validation gradient provides a stable direction of improvement. However, when the model is ill-trained or under-trained (e.g., far from convergence), gradients can be noisy or unstable, and the validation gradient may not yet be informative about the final optimization geometry. In such cases, the gradient-alignment signal can become unreliable, and IP may fail to rank samples correctly.
@@ 4 Correctness Verification on Synthetic Data
We evaluate the effectiveness of our Inner Product (IP) score as a surrogate for influence functions on two synthetic datasets representing convex and non-convex optimization settings.
For the convex case, we generate a linearly separable dataset using scikit-learn make_blobs function, consisting of 150 training samples and 100 test samples. We manually flip the labels of 10 training samples to introduce noise. A logistic regression model is trained on this dataset to assess IP under convex optimization. For the non-convex case, we use the half-moons dataset generated by make_moons, with 200 training samples (including 20 label-flipped noisy samples, 10 per class) and 100 test samples. We train a three-layer Multi-Layer Perceptron (MLP) with fully connected layers and ReLU activations. The network maps the input to 32 hidden units, maintains this dimensionality in the second layer, and outputs a single sigmoid-activated neuron for binary prediction.
Figure 1A–B visualize the linear dataset, while D–E present the half-moons dataset. Figures 1 C and F report the influence scores and IP scores for each training sample, computed using Eqs. (1) and (2). In the convex setting, the IP score closely matches the influence score, showing near-perfect correlation and order consistency. Detrimental samples receive negative scores under both measures, whereas beneficial or neutral samples yield positive or near-zero values. In contrast, in the non-convex setting, influence scores become less reliable: detrimental samples are intermingled with normal samples, likely due to inaccuracies in Hessian inverse approximation. The IP score, however, continues to clearly separate detrimental samples from inliers. As discussed previously, a low IP score reflects strong divergence between a training sample and the validation objective, indicating negative contribution. Notably, despite its simplicity, IP performs comparably to more complex methods that rely on Hessian-inverse approximations, even in the non-convex case. Additional empirical evaluations on model fairness and robustness are provided in Appendix A.1.
Next, we will present a unified empirical program along three sections: (1) noisy-label data curation for vision, where we rank samples by influence, and retrain to measure accuracy gains; (2) fairness-aware curation for NLP, where we compute influence on both utility and a demographic-perturbation fairness score, and visualize the accuracy-fairness frontier; (3) robustness to adaptive test-time evasion, where we compare pre/post-attack accuracy and evaluate three defenses.
5 Noisy Label Correction for Vision Datasets
In this section, we demonstrate the effectiveness of our IP Ensemble in identifying detrimental samples on noisy vision datasets. Specifically, we choose three benchmark datasets CIFAR-10N [61], CIFAR-100N [61], and Animal-10N [51] in the noisy label learning area. Both the CIFAR-10N and CIFAR-100N datasets [44] consist of the same input images as their clean counterparts, CIFAR-10 (10 classes) and CIFAR-100 (100 classes) [30], respectively. Each input is a 32x32 RGB image with dimensions (3,32,32). However, for CIFAR-10N and CIFAR-100N, the labels contain real-world human annotation errors collected using three annotators on Amazon Mechanical Turk. Since these datasets are based on human-annotated noise, they provide a more realistic modeling of noisy real-world scenarios compared to synthetic alternatives. The training set for both datasets contains 50,000 image-label pairs, while the test set contains 10,000 clean image-label pairs. Specifically, CIFAR-10N encompasses three distinct noise settings: aggregate (“-a”), random (“-r”), and worst (“-w”). The “aggregate” setting has the lowest noise rate (9.03%), where labels are determined via majority voting among three annotators, with ties resolved randomly. The “random” setting has intermediate noise (17.23%), adopting labels from the first annotator. The “worst” setting has the highest noise rate (40.21%), deliberately selecting the most inconsistent annotation per sample.
For competitive methods, we choose the following influence function-based methods. TracIn [47] replaces the Hessian matrix with the identity matrix and considers checkpoints during the training process; LiSSA [27] and EKFAC [20] employ implicit Hessian-vector products and Kronecker-Factored curvature to efficiently approximate the inverse of the Hessian matrix; DataInf [31] swaps the order of matrix multiplication for obtaining a closed-form expression; Self-TracIn [53] and Self-LiSSA [5] are the self-expression versions of TracIn and LiSSA, where the gradients of the validation set are replaced with , and only the last checkpoint, i.e., the converged model parameters, is used in Self-TracIn. GEX [26] utilizes ensemble methods based on checkpoints from extra stochastic gradient descent on the converged model. TDA [2] focuses on checkpoints during the training process, ensembling the influence via EKFAC [20]. Our IP Ensemble method constitutes an ensemble version of IP with dropout applied on model parameters and an ensemble size of 5.
TracIn, GEX, IP, and IP Ensemble are all Hessian-free ensemble methods except for IP. To clarify their differences, we present their calculations explicitly:
where in IP represents the converged model parameters; denotes the checkpoints saved during training; is derived by additional training on the converged model; IP Ensemble obtains through dropout mechanisms on the converged model; is the ensemble size. The advantages of IP Ensemble are that it does not require checkpoint saving as in TracIn, nor does it necessitate extra training iterations as in GEX. We argue that early checkpoints may not effectively reflect sample influence, and additional training on converged models has minimal impact. We provide our implementation in an anonymous open-source repository: https://anonymous.4open.science/r/IP_ensemble-BA0F/README.md. The experiments were conducted on a Linux (Ubuntu 20.04.6 LTS) server using NVIDIA GeForce RTX 4090 GPUs with 24GB VRAM running CUDA version 12.3 and driver version 545.23.08.